OpenStudy (anonymous):

Find the solution to the equation 64^(2 – x) = 4^4x.

OpenStudy (anonymous):

its a log function

jhonyy9 (jhonyy9):

64=4^4 so than you can writing like 4^4(2-x) =4^4x so than for the left part being equal with right part so than need to be those exponents equales so than 4(2-x)=4x 8-4x=4x 8=8x x=8/8 =1 x=1